OpenStudy (anonymous):

If F(x) = 5x - 3 and G(x) = x2 + 3, what is G(F(x)?

hero (hero):

Hint: G(x) = x^2 + 3 G(F(x)) = (F(x))^2 + 3 G(5x - 3) = (5x - 3)^2 + 3 Simplify it

hero (hero):

@stephanied2013, did you finish simplifying it?